Hi Can any one help me out with this prob.
There are three blue marbles, three red marbles , and three yellow marbles in a bowl. What is the probability of selecting exactly one red marble from the bowl after three successive marbles are withdrawn from the bowl?

Answer by stanbon(48569)

There are three blue marbles, three red marbles , and three yellow marbles in a bowl. What is the probability of selecting exactly one red marble from the bowl after three successive marbles are withdrawn from the bowl?
------------
P(red) = 3/9 = 1/3
P(not red) = 2/3
-------------------
P(exactly one red in three draws) = 3C1(1/3)(2/3)^2 = 3(1/3)(4/9) = 4/9
=================
Cheers,
Stan H.